Part 2 Round Robin Scheduling Process Burst Time in Ms P1 4 P2 2 P3 1...

Part 2

Round Robin Scheduling

Process

Burst Time in Ms

P1

4

P2

2

P3

1

Calculate the average waiting time and turnaround time using round robin scheduling, where the time quantum q = 2 ms

According to the Round Robin algorithm, the arrival of processes is shown in the following Gantt chart.

P1

P2

P3

P1

P2

P3

P1

P2

P3

P1

0                 2                 4                  6                8                 10               12                14               15              16               18

From the Gantt chart, the waiting time and the turnaround time are calculated as:

Waiting time for P1 =

Waiting time for P2 =

Waiting time for P2 =

Therefore, the average waiting time =

Turnaround time=
Answer #1

To find out average turnaround time, we need to calculate COMPLETION TIME.

Completion time for P1 = 18

Completion time for P2 = 15

Completion time for P3 = 16

Now, average turnaround time = 18 + 15 + 16/ 3

= 16.33

Turn Around Time: total time the process stays in system

Waiting time for P1 = 16 - 6 = 10

Waiting time for P2 = 14 - 4 = 10

Waiting time for P3 = 15 - 4 = 11

average waiting time = (10 + 10 + 11)/3 = 10.33

Waiting Time: total time the process needs to wait for there complete execution.